Find general solution of the differential equation


dy/dx=y2(1+ex)


1
Expert's answer
2021-06-01T04:39:23-0400

Let us find general solution of the differential equation "\frac{dy}{dx}=y^2(1+e^x)."


It is easy to see that "y=0" is a solution of this differential equation.


Assume that "y\ne 0". Then "\frac{dy}{y^2}=(1+e^x)dx".


It follows that "\int\frac{dy}{y^2}=\int(1+e^x)dx". Therefore, "-\frac{1}{y}=x+e^x+C."


We conclude that the solution is "y=0,\ \ -\frac{1}{y}=x+e^x+C."